OpenStudy (anonymous):

x^x = 16 How to solve this equation?

OpenStudy (bahrom7893):

take ln on both sides: Ln(x^x) = Ln(16)

OpenStudy (bahrom7893):

Using Ln properties: x*Ln(x) = 4*Ln(2)

OpenStudy (bahrom7893):

hmm that's ugly..

OpenStudy (bahrom7893):

sarkar probably has an idea.. cuz im out of them for now.

OpenStudy (bahrom7893):

false sarkar, 4^4 != 16

OpenStudy (bahrom7893):

neither is (-4)^(-4)

OpenStudy (bahrom7893):

sarkar, according to your answer, 16=256.... *facepalm*

OpenStudy (anonymous):

sorry for the error misinterpreted the question

OpenStudy (bahrom7893):

x^x means x to the power x, x times x means x*x

OpenStudy (bahrom7893):

4 times 4 is 16, 4 to the power of 4 is 256? anyone? or am I like totally wrong?

OpenStudy (anonymous):

you are correct bahrom7893

OpenStudy (anonymous):

i think result equals a dirty number.

OpenStudy (bahrom7893):

yay! okay, let's ask wolf what it thinks..

OpenStudy (bahrom7893):

http://www.wolframalpha.com/input/?i=Solve%28x^x%3D16%2Cx%29 So x = 2.7453...

OpenStudy (anonymous):

Yeah but how would you get that dirty number?

OpenStudy (bahrom7893):

good question ishaan, so u get a good answer

OpenStudy (anonymous):

nooo he is wrong no offense 4^4 is 256

OpenStudy (anonymous):

the answer is 2^4

OpenStudy (anonymous):

2^4=16 because when you have an exponent you have to multiply the base number (2) by the eponent(4) so it would be like this... 2x2x2x2= 16

OpenStudy (anonymous):

astrid are you nice :D is X is 2 or 4 ?

OpenStudy (anonymous):

I am convinced by Luis Rivera's reply

OpenStudy (anonymous):

alright well i gave it to you the easy way, also we have different ways of solving so that means you have a chance to choose, n no problem good luck

OpenStudy (anonymous):

astrid farnsworth i like you : )

OpenStudy (bahrom7893):

astrid, x^x is not 2^4.... *another facepalm*

OpenStudy (bahrom7893):

2 is not equal to 4.